# Rethinking the Starting Point: Collaborative Pre-Training for Federated Downstream Tasks

### v3

@@ Abstract
A few recent studies have demonstrated that leveraging centrally pre-trained models can offer advantageous initializations for federated learning (FL). However, existing pre-training methods do not generalize well when faced with an arbitrary set of downstream FL tasks. Specifically, they often (i) achieve limited average accuracy, particularly when there are unseen downstream labels, and (ii) result in significant accuracy variance, failing to provide a balanced performance across clients. To address these challenges, we propose CoPreFL, a collaborative/distributed pre-training approach which provides a robust initialization for downstream FL tasks. The key idea of CoPreFL is a model-agnostic meta-learning (MAML) procedure that tailors the global model to closely mimic heterogeneous and unseen FL scenarios, resulting in a pre-trained model that is rapidly adaptable to arbitrary FL tasks. Our MAML procedure incorporates performance variance into the meta-objective function, balancing performance across clients rather than solely optimizing for accuracy. Through extensive experiments, we demonstrate that CoPreFL obtains significant improvements in both average accuracy and variance across arbitrary downstream FL tasks with unseen/seen labels, compared with various pre-training baselines. We also show how CoPreFL is compatible with different well-known FL algorithms applied by the downstream tasks, enhancing performance in each case.
@@ 1 Introduction
Federated learning (FL) has emerged as a popular distributed machine learning paradigm, facilitating collaborative model training among sets of clients through periodic aggregations of local models by a server [31, 24]. In recent years, significant research attention has been given to various components of the FL process, such as aggregation schemes [21, 42] or local training techniques [39, 40]. One aspect that remains understudied, however, is the impact of model initialization in FL. While pre-training has demonstrated notable performance improvements in centralized AI/ML [36, 11, 14], most existing FL works initialize training using random weights, rather than starting from well pre-trained models.
@@
Motivation. Recently, a few works [34, 8] have shown that initializing FL with centrally pre-trained models can enhance the resulting average performance across clients. However, existing centralized pre-training methods face important drawbacks in practice, where they often have to handle newly emerging and/or heterogeneous downstream FL tasks unanticipated during pre-training. These drawbacks include: (i) limited average accuracy (despite outperforming random initialization), due to newly encountered data and labels, and (ii) large performance variance, failing to provide balanced accuracy across clients. The histograms in Figure 1, which compare the performance of different pre-trained models in various downstream image classification tasks (see Section 4 for details), illustrate these limitations. Compared to random initialization, although utilizing the centrally pre-trained model enhances average accuracy, it introduces substantial performance variance across clients, a well-cited concern in distributed AI/ML [29, 9]. Furthermore, we see that the achievable average accuracy of centralized pre-training is also suboptimal. This indicates that such models struggle to mimic data heterogeneity and other diverse characteristics present in downstream FL tasks.
@@
Figure 1: (Left): Overview of CoPreFL, aiming to provide a robust initialization for an arbitrary set of downstream FL tasks. (Right): Average accuracy and variance achieved by FL tasks (from Sec. 4) initialized by various pre-trained models. Centralized pre-training achieves limited performance as it is not able to capture the heterogeneous characteristics of unforseen FL settings. CoPreFL demonstrates improved performance in terms of both average accuracy and variance by strategically mimicing downstream FL scenarios during pre-training.
@@
Goals. Motivated by these limitations, the goal of this paper is to develop a robust FL pre-training methodology that provides a model initialization which achieves two main objectives: (i) improved average accuracy, and (ii) reduced performance variance to achieve balanced accuracy across clients, in each downstream FL task encountered. This is particularly challenging as it must be achieved across an arbitrary set of downstream FL tasks, with the possibility of facing data statistics and labels that were unseen during pre-training, due to time-varying environments (e.g., self-driving cars confronting previously unseen objects to classify), new clients joining the system (e.g., face/speech recognition for new phone users), or other factors. Therefore, the pre-trained model must adapt well to unfamiliar classes and accommodate new data heterogeneity during downstream FL tasks, a challenge overlooked by existing methods [34, 8]. We summarize our research question as follows:
@@
  • Distributed pre-training infused with meta-learning: The foundation of CoPreFL is an FL-inspired pre-training procedure which iteratively applies model-agnostic meta-learning (MAML)-based updates on the collaboratively-constructed global model. Through our developed MAML procedure, CoPreFL produces robust initializations in the sense of enabling the pre-trained model to easily adapt to unseen labels and heterogeneous data distributions encountered in an arbitrary set of downstream tasks. We note that this approach differs in purpose and method from prior works that use meta-learning for personalization in FL [7, 22, 15]: since our downstream tasks aim to construct a global model rather than client-specific personalized models, we conduct meta-updates based on the global model instead of directly using local models. The limitations of existing meta-learning methods for pre-training will be empirically demonstrated in Section 4.2.
  • Meta-objective function incorporating variance: To enhance average accuracy while improving performance balance among clients in the downstream FL tasks, we explicitly incorporate both expected loss and performance variance into the meta-objective function employed during pre-training in CoPreFL. In doing so, we introduce a first-order approximation for efficiently computing the gradient of the proposed meta-objective function.
  • Relaxing the assumption of centrally stored pre-training data: CoPreFL relaxes the assumption made by existing works that all pre-training data is stored centrally. Specifically, we develop our pre-training algorithm for a range of hybrid client-server data storage settings, where (i) data is exclusively held by distributed clients, or (ii) the server also maintains a portion of data. This aligns with many practical distributed/FL applications that possess data privacy limitations and/or communication constraints. Our approach is applicable and advantageous even when all pre-training data is centrally stored, which is also validated through our experiments.
  • Extensive experiments across various downstream FL tasks and settings: We compare CoPreFL with various pre-training baselines, across diverse downstream FL tasks, with several different combinations of data distributions and seen/unseen labels, and different client-server pre-training data allocations. We observe significant improvements in both average accuracy and performance variance achieved across the clients in each case where the downstream tasks are initialized based on CoPreFL. We also show how CoPreFL is compatible with different popular FL variants employed for downstream training. Finally, we demonstrate that CoPreFL’s improvements extend to distributional shifts between pre-training and downstream tasks.
Our work is among the first to consider FL in both the pre-training and downstream stages of distributed learning tasks. In doing so, we introduce several unique features tailored to FL, including meta-updating the global model during distributed pre-training, hybrid client-server learning, and balancing between average performance and variance across the clients.
@@ 2 Related Work
Pre-training for FL. While pre-training has been extensively studied in centralized AI/ML applications [36, 4, 11, 14], its effects on downstream FL tasks have remained relatively unexplored. A few recent works have studied the effect of model initialization in FL [34, 8], showing that conducting FL initialized from centrally pre-trained models can significantly enhance average performance compared with random initializations. However, as observed in Figure 1, such initialization strategies introduce large performance variance and even limited average accuracy, since they are not able to mimic multiple downstream FL settings. To overcome this, we develop a MAML-based pre-training strategy tailored to distributed downstream settings, so that initializing FL with the pre-trained model can improve both average accuracy and performance variance while addressing the challenges of heterogeneous/unseen data encountered in each downstream FL task. We demonstrate the advantages of CoPreFL over existing pre-training schemes through extensive experiments in Section 4.
Meta-learning in FL. CoPreFL employs meta-learning, aiming to construct a global model that is adaptable to arbitrary downstream FL tasks while jointly optimizing performance and variance. This distinguishes it from other FL methods that have employed meta-learning, such as personalized FL [7, 22, 15] and few-round FL [35]. Specifically, [35] employs a meta-learning-based episodic training strategy to adapt FL to any client group within only a few rounds of FL. However, [35] does not consider performance imbalance across clients as an objective. Further, the practical scenario in which the server holds a proxy dataset for hybrid pre-training is not considered. In personalized FL [7, 22, 15], meta-update is conducted based on the performance of each client’s local model, aiming for better personalization to clients in downstream tasks. By contrast, our emphasis is on establishing a pre-trained model that results in high-accuracy and equitable global models for downstream tasks that themselves are trained through FL, requiring meta-updates based on the performance of the global model. We will see in Section 4 that CoPreFL leads to significant performance improvements over existing meta-learning strategies.
Performance imbalance in FL. Several works in the FL literature have considered performance imbalance [32, 29, 9], typically aiming to construct a global model that satisfies as many clients as possible (e.g., achieving a uniform distribution of accuracy across participants). When properly architected, such models have a higher likelihood of satisfying new clients joining the FL system without requiring additional model training. These methods can be applied downstream from our pre-training methodology, whose primary objective is to construct a robust initial model that will lead to higher average and more balanced performance across clients after FL training. Our experiments in Section 4 will show how initializing these methods from CoPreFL further enhances their performance.
@@ 3.1 Problem Setup and Pre-Training Objectives
Federated downstream tasks. Referring to Figure 1, CoPreFL aims to provide a robust initialization for an arbitrary set of downstream tasks, each governed by FL. In each of these downstream tasks, we assume a central server/aggregator is connected to a set of clients $G$. Starting from the initialized model $w^{0}$, each FL task iterates between (i) parallel local training at the clients and (ii) global aggregations at the server, across multiple communication rounds. Specifically, in each downstream training round $r$, every client $g\in G$ downloads the previous global model $w^{r-1}$ from the server and subsequently updates it through multiple iterations of stochastic gradient descent (SGD) using its local dataset, denoted $D_{g}$. After completing their local updates, clients upload their updated models, denoted $w^{r}_{g}$, to the server for aggregation. This aggregation results in a new global model $w^{r}=\sum_{g\in G}\frac{|D_{g}|}{|D|}w^{r}_{g}$ assuming FedAvg [31] is employed, where $|D|$ is the total data samples across all clients. This entire process is iterated over $r=1,...,R$ communication rounds for each task.
@@
Scenario I emulates the setting encountered by the downstream FL tasks for their training. A key challenge is that the labels and data that appear in the pre-training stage may be different from the ones that will appear in the downstream tasks. Scenario II emulates settings where the server providing the downstream model initializations holds some data that reflects the broader population distribution (e.g., a self-driving car manufacturer with a data store of images encountered on roadways). Such hybrid FL settings that combine client data with a relatively small portion of server data are becoming popular [43, 3], but have not been investigated for pre-training. Further, as we will discuss in Remark 2, our method is still applicable even in settings where all pre-training data is centralized.
Pre-training objectives. Our goal is to design a pre-trained model $\Phi^{*}$ that serves as a robust starting point $w^{0}=\Phi^{*}$ for any arbitrary downstream FL task, as shown in Figure 1. More precisely, letting $\mathcal{G}$ be a random variable representing possible client sets comprising a downstream FL task, one of the goals in designing $\Phi^{*}$ is to minimize the following objective function:
@@
where $p(\mathcal{G})$ represents the probability distribution over $\mathcal{G}$, $G$ is a specific client group (i.e., a specific task) drawn from $p(\mathcal{G})$, $f(\cdot)$ is the per-client loss function for downstream training, $w^{R}(\Phi,G)$ symbolizes the final $R$-th round global model derived from client set $G$ when initialized by $\Phi$, and $D_{g}$ represents the local dataset of client $g$. $A(\Phi)$ represents the average FL performance across all clients comprising downstream tasks, with each group weighted by likelihood of occurrence.
On the other hand, FL settings can lead to significant performance variations among clients, especially when the aggregated models are biased towards those with larger datasets. This performance variation can be assessed by quantifying the variance in testing accuracy across participants [29]. Therefore, besides achieving performance gains on any unseen FL task, we also aim for the final global model $w^{R}(\Phi^{*},G)$ initialized from our designed pre-trained model $\Phi^{*}$ to exhibit a balanced testing performance distribution across the client set $G$. Specifically, our second objective in designing $\Phi^{*}$ is to minimize the variance of the loss distribution across participants in downstream FL tasks, i.e.,
$F(\Phi)=\mathbb{E}_{{G}\sim p(\mathcal{G})}\biggl{[}\frac{1}{|G|}\sum_{g\in{G}}f^{2}(w^{R}(\Phi,G),D_{g})-\Big{(}\frac{1}{|G|}\sum_{g\in{G}}f(w^{R}(\Phi,G),D_{g})\Big{)}^{2}\biggr{]}.$ (2)
Overview of approach. One of our key contributions is striking a balance between (1) and (2). This is challenging as $D_{g}$, $\mathcal{G}$, and $p(\mathcal{G})$ are not known during pre-training, preventing us from directly optimizing $A(\Phi)$ and $F(\Phi)$. To address this, we develop a model-agnostic meta-learning (MAML) approach for CoPreFL to mimic statistical heterogeneity found in downstream FL. This approach yields pre-trained models that offer robust initialization for unseen downstream tasks, considering (1) and (2) in the MAML formulation. Specifically, in Sections 3.2 and 3.3, we construct a pre-training environment for scenarios I and II mirroring arbitrary federated setups encountered downstream, facilitating the pre-trained model’s ability to account for data heterogeneity across clients and tasks. Our meta-learning-based CoPreFL updates the pre-trained model iteratively over federated rounds using a support set, followed by a concluding adjustment (meta-update) using a query set. By treating the query set as unseen knowledge, our pre-trained model has the capability to effectively handle unforeseen FL scenarios in downstream tasks while striking a balance between (1) and (2).
@@ 3.2 CoPreFL in Scenario I (Pre-training with Distributed Clients)
3:     Randomly select a set of clients $m\subset M$ to participate in learning
@@
8:           $\ell^{e}_{S_{j}}(\Phi^{t})\leftarrow\frac{1}{|S_{j}|}\sum_{(x,y)\in S_{j}}\ell(\Phi^{t,e}_{j}(x),y)$// Compute local support loss at each epoch}
9:           $\Phi^{t,e}_{j}$leftarrow $\Phi^{t,e-1}_{j}$- $\eta\nabla\ell^{e}_{S_{j}}(\Phi^{t})$// Perform SGD local update using support loss}
@@
12:     $\overline{\Phi^{t}}$leftarrow $\sum_{j\in m}\frac{|S_{j}|}{\sum_{i\in m}|S_{i}|}\Phi^{t,E}_{j}$// Model aggregation to construct temporary global model}
@@
15:        $\mathcal{L}_{Q_{j}}(\overline{\Phi^{t}})$leftarrow $\frac{1}{|Q_{j}|}$ $\sum_{(x,y)\in Q_{j}}{}\ell(\overline{\Phi^{t}}(x),y)$// Compute local loss (and gradient) using query set $Q_{j}$}
@@
18:     $\mathcal{L}_{meta}(\overline{\Phi^{t}})=\gamma\mathcal{L}_{Q}(\overline{\Phi^{t}})+(1-\gamma)\sigma_{Q}^{2}(\overline{\Phi^{t}})$// Customized query meta-loss}
19:     $\Phi^{t}$leftarrow $\overline{\Phi^{t}}$- $\zeta\nabla\mathcal{L}_{meta}(\overline{\Phi^{t}})$// Meta-learning model update using customized loss}
@@
We first consider the scenario where pre-training data is contained at a set of $M$ distributed clients, and no data is stored on the server. The detailed procedure of CoPreFL for this case is given in Algorithm 1. To start, in each round $t=1,...,T$ of pre-training, a set of clients $m\subset M$ is randomly selected to participate in the current round. Further, each participating client $j\in m$ splits its local pre-training dataset $D^{p}_{j}=S_{j}\cup Q_{j}$ into a support set $S_{j}$ and query set $Q_{j}$, which are disjoint. These steps emulate downstream task variations within and across each pre-training round (i.e., by changing the participating clients across rounds, and holding out the query sets in each round), allowing our meta-learning to maximize the model’s generalization ability in unseen downstream scenarios.
Temporary pre-training model construction. In each round $t$, participating clients $j\in m$ download $\Phi^{t-1}$ from the server. Subsequently, clients undertake a series of local training epochs using their respective support sets $S_{j}$, resulting in a local support loss $\ell^{e}_{S_{j}}(\Phi^{t})$ for each epoch $e$, defined in line 8 of Algorithm 1, where $\ell(\cdot)$ denotes the per-datum loss function (e.g., cross-entropy loss for classification). After all participants finish $E$ epochs, we obtain the updated local model $\Phi^{t,E}_{j}$. Clients then send their updated models to the server for aggregation, resulting in $\overline{\Phi^{t}}$ (defined in line 12). This model can be viewed as the temporary pre-training model that will be further refined by our methodology using the query sets, with the objective of obtaining robust global models at the conclusion of downstream task training.
@@
$\mathcal{L}_{Q}(\overline{\Phi^{t}})=\sum_{j\in m}\mathcal{L}_{Q_{j}}(\overline{\Phi^{t}})\;\;\;\;\text{ and}\;\;\;\;\sigma_{Q}^{2}(\overline{\Phi^{t}})=\frac{1}{|m|}\sum_{j\in m}\Big{(}\mathcal{L}_{Q_{j}}(\overline{\Phi^{t}})-\frac{1}{|m|}\mathcal{L}_{Q}(\overline{\Phi^{t}})\Big{)}^{2},$ (3)
where $\mathcal{L}_{Q_{j}}$ represents the loss evaluated using query sets $Q_{j}$ of participants, $\mathcal{L}_{Q}$ denotes the overall query loss (characterized by aggregating $\mathcal{L}_{Q_{j}}$ across all participants), and $\sigma_{Q}^{2}$ represents the performance variance evaluated using the query set across participants. Based on this, to balance the performance-variance tradeoff, we construct a customized query meta-loss function $\mathcal{L}_{meta}(\overline{\Phi^{t}})$ to minimize not only the overall query loss $\mathcal{L}_{Q}(\overline{\Phi^{t}})$ when encountering unseen data, but also the variance $\sigma_{Q}^{2}(\overline{\Phi^{t}})$ of query losses across participants. Formally, we aim to solve
@@
where $\gamma\in[0,1]$ represents a controllable balancer between the average performance and variance. Setting $\gamma=0$ encourages a more uniform/balanced training accuracy distribution, aligning with $\sigma_{Q}^{2}$, but may sacrifice average performance. A larger $\gamma$ emphasizes the average performance with less consideration for uniformity, optimizing the pre-trained model more towards $\mathcal{L}_{Q}$.

### v4

@@ Abstract
A few recent studies have shown the benefits of using centrally pre-trained models for initializing federated learning (FL). However, existing pre-training methods do not generalize well when faced with an arbitrary set of downstream FL tasks. Specifically, they often (i) achieve limited average accuracy, particularly when there are unseen downstream labels, and (ii) result in significant accuracy variance, failing to provide a balanced performance across clients. To address these challenges, we propose CoPreFL, a collaborative/distributed pre-training approach which robustly initializes for downstream FL tasks. The key idea of CoPreFL is a model-agnostic meta-learning (MAML) procedure that tailors the global model to closely mimic heterogeneous and unseen FL scenarios, resulting in a pre-trained model that is rapidly adaptable to any FL task. Our MAML procedure integrates performance variance into the meta-objective function, balancing performance across clients rather than solely optimizing for accuracy. Extensive experiments show that CoPreFL significantly enhances both average accuracy and reduces variance in arbitrary downstream FL tasks with unseen/seen labels, outperforming various pre-training baselines. Additionally, CoPreFL proves compatible with different well-known FL algorithms applied by the downstream tasks, boosting performance in each case.
@@ 1 Introduction
Federated learning (FL) has gained prominence as a distributed machine learning framework, enabling collaborative training among clients by periodic aggregations of local models on a server (McMahan et al. 2017; Konecný et al. 2016). Recent research has extensively explored various aspects of FL, such as aggregation schemes (Ji et al. 2019; Wang et al. 2020) or local training techniques (Reddi et al. 2021; Sahu et al. 2018). One aspect that remains understudied, however, is the impact of model initialization in FL. While pre-training boosts performance in centralized AI/ML (Radford et al. 2019; Devlin et al. 2019; Dosovitskiy et al. 2021), most FL works still rely on random weight initialization instead of well pre-trained models.
@@
Motivation. Recently, a few works (Nguyen et al. 2023; Chen et al. 2023) have shown that initializing FL with centrally pre-trained models can enhance the average performance across clients. Yet, existing centralized pre-training methods face significant challenges, particularly when handling newly emerging and/or heterogeneous downstream FL tasks unanticipated during pre-training. These include: (i) limited average accuracy (despite outperforming random initialization), due to newly encountered data and labels, and (ii) large performance variance, resulting in unbalanced accuracy across clients. The histograms in Figure 1 show the performance of various pre-trained models in multiple downstream image classification tasks (see Section 4 for details), illustrating these limitations. While using centrally pre-trained models improves average accuracy over random initialization, it introduces significant performance variance across clients, a well-cited concern in distributed AI/ML (Li et al. 2020; Cho et al. 2022). Additionally, the achievable average accuracy of centralized pre-training remains suboptimal, indicating that such models struggle to mimic data heterogeneity and diversity in downstream FL tasks.
@@
Figure 1: (Left): Overview of CoPreFL, aiming to provide a robust initialization for an arbitrary set of downstream FL tasks. (Right): Average accuracy and variance achieved by FL tasks (from Section 4) initialized by various pre-trained models. Centralized pre-training achieves limited performance as it is not able to capture the heterogeneous characteristics of unforseen FL settings. CoPreFL demonstrates improved performance in terms of both average accuracy and variance by strategically mimicing downstream FL scenarios during pre-training.
@@
Goals. Motivated by these limitations, we aim to develop a robust FL pre-training methodology that provides an initialization which achieves two main objectives: (i) improved average accuracy, and (ii) reduced performance variance for balanced accuracy across clients, in each downstream FL task. This is challenging as it must work across any arbitrary set of downstream FL tasks, which may include data and labels previously unseen due to factors like time-varying environments or new clients joining the system. Thus, the pre-trained model must handle unfamiliar classes and data heterogeneity during downstream FL tasks, a challenge overlooked by existing methods (Nguyen et al. 2023; Chen et al. 2023). We summarize our research question as follows:
@@
  • Distributed pre-training infused with meta-learning: The core of CoPreFL is an FL-inspired pre-training procedure which employs model-agnostic meta-learning (MAML)-based updates on the collaboratively-bulit global model. Through our developed MAML procedure, CoPreFL ensures robust initializations, enabling the pre-trained model to easily adapt to unseen labels and various data distributions in any downstream FL tasks. This approach differs in purpose and method from prior works that use meta-learning for personalization in FL (Chen et al. 2018; Jiang et al. 2019; Fallah, Mokhtari, and Ozdaglar 2020): since our downstream tasks aim to construct a global model rather than client-specific personalized models, we conduct meta-updates based on the global model instead of directly using local models.
  • Meta-objective function incorporating variance: To enhance average accuracy while improving performance balance among clients in the downstream FL tasks, we explicitly incorporate both expected loss and performance variance into the meta-objective function during pre-training in CoPreFL. In doing so, we introduce a first-order approximation for efficiently computing the gradient of the proposed meta-objective function.
  • Relaxing the assumption of centrally stored pre-training data: CoPreFL relaxes the assumption made by existing works that all pre-training data is stored centrally. We develop our pre-training algorithm for various hybrid client-server data storage settings, where (i) data is exclusively held by distributed clients, or (ii) the server also maintains partial data. This is crucial for distributed/FL applications with data privacy limitations. Our approach is applicable even when all pre-training data is centrally stored, as confirmed by our experiments.
  • Extensive experiments across various downstream FL tasks and settings: We compare CoPreFL against various baselines across diverse downstream FL tasks with different data distributions, seen/unseen labels, and client-server data allocations. Results show notable improvements in accuracy and performance variance when downstream tasks are initialized with CoPreFL. We also show CoPreFL’s compatibility with various popular FL algorithms used downstream and its resilience to distributional shifts.
Our work is among the first to consider FL in both the pre-training and downstream stages of distributed learning tasks. We introduce several unique features tailored to FL, including meta-updating the global model during distributed pre-training, hybrid client-server learning, and balancing between average performance and variance across the clients.
@@ 2 Related Work
Pre-training for FL. While pre-training has been extensively studied in centralized AI/ML applications (Radford et al. 2019; Brown et al. 2020; Devlin et al. 2019; Dosovitskiy et al. 2021), its effects on downstream FL tasks remains less explored. A few recent works have shown that starting FL with centrally pre-trained models can improve performance over random initializations (Nguyen et al. 2023; Chen et al. 2023). However, as observed in Figure 1, such initialization strategies often lead to large performance variance and even limited average accuracy, since they are not able to mimic multiple downstream FL settings. To address this, we develop a MAML-based pre-training strategy, CoPreFL, tailored for distributed downstream settings, so that initializing FL with this pre-trained model can improve both average accuracy and performance variance while addressing the challenges of heterogeneous/unseen data encountered in downstream FL tasks.
Meta-learning in FL. CoPreFL utilizes meta-learning to construct a global model adaptable to any downstream FL tasks, optimizing both performance and variance. This distinguishes it from other meta-learning-based FL, like personalized FL (Chen et al. 2018; Jiang et al. 2019; Fallah, Mokhtari, and Ozdaglar 2020; Chu et al. 2022) and few-round FL (Park et al. 2021). Specifically, Park et al. (2021) uses meta-learning-based episodic training for quick FL adaptation, but it overlooks performance imbalance across clients and fails to consider scenarios where the server holds a proxy dataset for hybrid pre-training. In personalized FL, meta-update target individual client model performance, aiming for better personalization to clients in downstream tasks. In contrast, our focus is on developing a pre-trained model that ensures high accuracy and fairness for global models in downstream FL tasks.
Performance imbalance in FL. Several works in FL address performance imbalance (Mohri, Sivek, and Suresh 2019; Li et al. 2020; Cho et al. 2022) typically by creating a global model that satisfies as many clients as possible (e.g., achieving a uniform accuracy across clients). Such models are more likely to satisfy new clients joining the FL system. These methods can be applied downstream from our pre-training methodology, whose primary objective is to build a robust initial model that will lead to higher average and more balanced performance across clients after FL training.
@@ 3.1 Problem Setup and Pre-Training Objectives
Federated downstream tasks. Referring to Figure 1, CoPreFL aims to provide a robust initialization for any downstream FL tasks. In each of these downstream tasks, we assume a central server is connected to a set of clients $G$. Starting from the initialized model $w^{0}$, each FL task iterates between (i) local training at the clients and (ii) global aggregations at the server, across multiple communication rounds. In each downstream round $r$, every client $g\in G$ downloads the previous global model $w^{r-1}$ from the server and subsequently updates it through multiple iterations of stochastic gradient descent (SGD) using its local dataset, denoted $D_{g}$. After completing their local updates, clients upload their updated models, denoted $w^{r}_{g}$, to the server for aggregation. This aggregation results in a new global model $w^{r}=\sum_{g\in G}\frac{|D_{g}|}{|D|}w^{r}_{g}$ assuming FedAvg (McMahan et al. 2017) is employed, where $|D|$ is the total data samples across all clients. This entire process is iterated over $r=1,...,R$ communication rounds for each task.
@@
Scenario I simulates downstream FL tasks where pre-training labels and data may differ from those in downstream tasks. Scenario II represents settings where the server holds data reflecting the broader population distribution (e.g., a self-driving car manufacturer with database of roadway images). Such hybrid FL settings that combine client data with a relatively small portion of server data are becoming popular (Yang, Chen, and Shen 2023; Bian et al. 2023), but remain underexplored for pre-training. Further, as we will discuss in Remark 2, our method is still applicable even when all pre-training data is centralized.
Pre-training objectives. Our goal is to design a pre-trained model $\Phi^{*}$ that serves as a robust starting point $w^{0}=\Phi^{*}$ for any arbitrary downstream FL task. More precisely, letting $\mathcal{G}$ represents possible client sets comprising a downstream FL task, the goal of designing $\Phi^{*}$ is to minimize the following objective function:
@@
where $p(\mathcal{G})$ represents the probability distribution over $\mathcal{G}$, $G$ is a specific client group (i.e., a specific task) drawn from $p(\mathcal{G})$, $f(\cdot)$ is the per-client loss function for downstream training, $w^{R}(\Phi,G)$ symbolizes the final $R$-th round global model derived from client set $G$ when initialized by $\Phi$, and $D_{g}$ represents the local dataset of client $g$. $A(\Phi)$ denotes the average FL performance across all clients fir downstream tasks, with each group weighted by likelihood of occurrence.
On the other hand, FL settings can lead to significant performance variations among clients, especially when the aggregated models are biased towards those with larger datasets. This performance variation can be measured by the variance in testing accuracy across participants (Li et al. 2020). Thus, besides improving performance for any FL task, we aim for the final global model $w^{R}(\Phi^{*},G)$ initialized from our pre-trained model $\Phi^{*}$ to achieve balanced testing performance across client set $G$. Specifically, our second objective for $\Phi^{*}$ is to minimize the variance of the loss distribution across participants in downstream FL tasks, i.e.,
$\begin{split}F(\Phi)=\mathbb{E}_{G\sim p(\mathcal{G})}&\biggl{[}\frac{1}{|G|}\sum_{g\in G}f^{2}(w^{R}(\Phi^{*},G),D_{g})\\ &-\biggl{(}\frac{1}{|G|}\sum_{g\in G}f(w^{R}(\Phi^{*},G),D_{g})\biggr{)}^{2}\biggr{]}.\end{split}$ (2)
Overview of approach. One of our key contributions is balancing (1) and (2). The challenge arises as $D_{g}$, $\mathcal{G}$, and $p(\mathcal{G})$ are unknown during pre-training, preventing us from directly optimizing $A(\Phi)$ and $F(\Phi)$. To address this, we develop a model-agnostic meta-learning (MAML) approach in CoPreFL to mimic statistical heterogeneity of downstream FL tasks. This method yields pre-trained models that offer robust initialization for unseen downstream tasks, considering (1) and (2). Detailed in Sections 3.2 and 3.3, we construct a pre-training environment for scenarios I and II that mirrors downstream federated setups, enabling the pre-trained model to handle data heterogeneity across clients and tasks. Our meta-learning-based CoPreFL updates the pre-trained model iteratively over federated rounds using a support set, with a concluding adjustment (meta-update) using a query set treated as unseen knowledge. This enables our pre-trained model to the effectively handle unforeseen FL scenarios downstream while balancing between (1) and (2).
@@ 3.2 CoPreFL in Scenario I (Pre-training with Distributed Clients)
3:     Randomly select a set of clients $m\subset M$ to participate
@@
8:           $\ell^{e}_{S_{j}}(\Phi^{t})\leftarrow\frac{1}{|S_{j}|}\sum_{(x,y)\in S_{j}}\ell(\Phi^{t,e}_{j}(x),y)$ Compute local support loss at each epoch}
9:           $\Phi^{t,e}_{j}$leftarrow $\Phi^{t,e-1}_{j}$- $\eta\nabla\ell^{e}_{S_{j}}(\Phi^{t})$ Perform SGD local update using support loss}
@@
12:     $\overline{\Phi^{t}}$leftarrow $\sum_{j\in m}\frac{|S_{j}|}{\sum_{i\in m}|S_{i}|}\Phi^{t,E}_{j}$ Model aggregation to construct temporary global model}
@@
15:        $\mathcal{L}_{Q_{j}}(\overline{\Phi^{t}})$leftarrow $\frac{1}{|Q_{j}|}$ $\sum_{(x,y)\in Q_{j}}{}\ell(\overline{\Phi^{t}}(x),y)$ Compute local loss (and gradient) using query set $Q_{j}$}
@@
18:     $\mathcal{L}_{meta}(\overline{\Phi^{t}})=\gamma\mathcal{L}_{Q}(\overline{\Phi^{t}})+(1-\gamma)\sigma_{Q}^{2}(\overline{\Phi^{t}})$ Customized query meta-loss}
19:     $\Phi^{t}$leftarrow $\overline{\Phi^{t}}$- $\zeta\nabla\mathcal{L}_{meta}(\overline{\Phi^{t}})$ Meta-learning model update using customized loss}
@@
We first consider the scenario where pre-training data is distributed across $M$ distributed clients, and no data is stored on the server. The detailed procedure of CoPreFL for this case is given in Algorithm 1. To start, in each round $t=1,...,T$ of pre-training, a set of clients $m\subset M$ is randomly selected to participate in the current round. Further, each participating client $j\in m$ splits its local pre-training dataset $D^{p}_{j}$ into a support set $S_{j}$ and query set $Q_{j}$, which are disjoint. These steps mimic downstream task variations within and across each pre-training round (i.e., by changing the participating clients across rounds, and holding out the query sets in each round), allowing our meta-learning to improve generalization in unseen downstream scenarios.
Temporary pre-training model construction. In each round $t$, participating clients $j\in m$ download $\Phi^{t-1}$ from the server. Subsequently, clients perform local training using their support sets $S_{j}$, yielding a local support loss $\ell^{e}_{S_{j}}(\Phi^{t})$ per epoch $e$, defined in line 8 of Algorithm 1, where $\ell(\cdot)$ is the per-datum loss function (e.g., cross-entropy for classification). After all participants finish $E$ epochs, we obtain the updated local model $\Phi^{t,E}_{j}$. Clients then send their updated models to the server for aggregation, resulting in $\overline{\Phi^{t}}$ (defined in line 12). This model can be viewed as the temporary pre-training model that will be further refined using query sets, with the objective of obtaining robust global models at the conclusion of downstream task training.
@@
$\begin{split}&\mathcal{L}_{Q}(\overline{\Phi^{t}})=\sum_{j\in m}\mathcal{L}_{Q_{j}}(\overline{\Phi^{t}})\text{ and}\\ &\sigma_{Q}^{2}(\overline{\Phi^{t}})=\frac{1}{|m|}\sum_{j\in m}\Big{(}\mathcal{L}_{Q_{j}}(\overline{\Phi^{t}})-\frac{1}{|m|}\mathcal{L}_{Q}(\overline{\Phi^{t}})\Big{)}^{2},\end{split}$ (3)
where $\mathcal{L}_{Q_{j}}$ represents the loss evaluated using query sets $Q_{j}$ of participants, $\mathcal{L}_{Q}$ denotes the overall query loss (characterized by aggregating $\mathcal{L}_{Q_{j}}$ across all participants), and $\sigma_{Q}^{2}$ represents the performance variance evaluated using clients’ query losses. To balance the performance-variance tradeoff, we construct a customized query meta-loss function $\mathcal{L}_{meta}(\overline{\Phi^{t}})$ to minimize both the overall query loss $\mathcal{L}_{Q}(\overline{\Phi^{t}})$ when encountering unseen data and the variance $\sigma_{Q}^{2}(\overline{\Phi^{t}})$ of query losses across participants. Formally, we aim to solve:
@@
where $\gamma\in[0,1]$ represents a balancer between the average performance and variance. Setting $\gamma=0$ encourages a more uniform accuracy distribution, aligning with $\sigma_{Q}^{2}$, but may sacrifice average performance. A larger $\gamma$ emphasizes the average performance with less consideration for uniformity, optimizing the pre-trained model more towards $\mathcal{L}_{Q}$.